OpenStudy (anonymous):

What is the directrix of the parabola 3y^2 - 12x = 0

OpenStudy (anonymous):

y^2 = 4x

OpenStudy (anonymous):

is the parabola..dont get excited:)

OpenStudy (anonymous):

see if a parabola is given b y^2 = 4ax its directrix is x+ a =0

OpenStudy (anonymous):

now in ur case a = 1 so x+1 = 0

OpenStudy (anonymous):

im so confused